The width of a rectangle is 3 less than twice its length. If the area of the rectangle is 33
cm^2
, what is the length of the diagonal?

Found 2 solutions by Boreal, Alan3354:
Answer by Boreal(15235)

length=x
width is 2x-3
area is 33 so 2x^2-3x=33
2x^2-3x-33=0
x=(1/4) (3+/- sqrt (9+264)); sqrt (273)=16.52
only positive root is x=4.88 cm
so length is 4.88 cm and width is 6.76 cm. Their product is 32.99 cm^2
the diagonal is the sqrt (4.88^2+6.76^2)=8.34 cm

Answer by Alan3354(69443)

The width of a rectangle is 3 less than twice its length. If the area of the rectangle is 33 cm^2
====================
L*W = 33
L = 2W - 3
-----
W*(2W-3) = 33
---------
Solve for W, then find L
Then find the diagonal.